# LLM-Based Multi-Hop Question Answering with Knowledge Graph Integration in Evolving Environments

### v1

@@ A.4 Prompt Setup and Post-Processing
The prompts used for edited fact triple extraction, relation chain extraction, and LLM-based QA are depicted in Figures  3,  4, and  5. The edited triple can be regarded as a specialized relation chain, with only one relation between entities and all entities known. All samples in the prompt are selected from the complete MQuAKE-CF dataset, ensuring they are distinct from the test samples.
@@
To improve the performance of LLMs in extracting relation chains and ensure that outputs conform to a specified format, we employ a 4-shot learning approach for the MQuAKE-CF dataset and a 3-shot learning approach for the MQuAKE-T dataset. For MQuAKE-CF, the approach involves presenting the model with samples of one 2-hop question, one 3-hop question, and two 4-hop questions. In contrast, for MQuAKE-T, the model is presented with one 2-hop question, one 3-hop question, and one 4-hop question.

### v2

@@ A.4 Prompt Setup and Post-Processing
The prompts used for edited fact triple extraction, relation chain extraction, and LLM-based QA are depicted in Figures 3, 4, and 5. The edited triple can be regarded as a specialized relation chain, with only one relation between entities and all entities known. All samples in the prompt are selected from the complete MQuAKE-CF dataset, ensuring they are distinct from the test samples.
@@
To improve the performance of LLMs in extracting relation chains and ensure that outputs conform to a specified format, we employ a 4-shot learning approach for the MQuAKE-CF dataset and a 3-shot learning approach for the MQuAKE-T dataset. For MQuAKE-CF, the approach involves presenting the model with samples of one 2-hop question, one 3-hop question, and two 4-hop questions. For MQuAKE-T, the model is presented with one 2-hop question, one 3-hop question, and one 4-hop question.